OpenStudy (kiakat):

HELP Prove: cos θ - cosθ •sin^2 θ = cos^3 θ. You must show all work. I understand how to do this problem kind of..but ive looked up a ton of different things and they simplified it to cosθ(1-sin^2θ)=cos^3θ How did they simplify that???

OpenStudy (kiakat):

@phi @amistre64

OpenStudy (kiakat):

those are all exponents in the equation btw

OpenStudy (amistre64):

post it correctly notated

OpenStudy (kiakat):

think i fixed it..

OpenStudy (amistre64):

a - ab = c how would you simplify it?

OpenStudy (kiakat):

distributive property?

OpenStudy (amistre64):

and what is our basic trig identity? the pythagorean one

OpenStudy (kiakat):

sin^2θ+cos^2θ=1

OpenStudy (amistre64):

well, an undistributive property so yeah a(1-b) = c good, so 1-sin^2 = what?

OpenStudy (kiakat):

cos^θ

OpenStudy (kiakat):

cos^2θ sorry

OpenStudy (amistre64):

then we are simplified show me the work

OpenStudy (kiakat):

using a(1-b)=c all you really have to do is place the values for a, b and c into the equation, right? So then it'd come out as cosθ(1-sin^2θ)=cos^3θ...which can be simplified into cosθ(cos^2θ) which is equal to cos^3θ

OpenStudy (amistre64):

very good

OpenStudy (kiakat):

thanks so much!!